# Quantifying the role of supernatural entities and the effect of missing data in Irish sagas

### v1

@@ 3.1 Core network properties
1 | Cúchulainn (70) | Cúchulainn (30) | Cúchulainn | Manannán |
2 | Fionn (69) | Fionn (19) | Manannán | Cúchulainn |
3 | Conchobhar (58) | Lugh (17) | Dagda | Dagda |
4 | Medb (41) | Naoise (16) | Lugh | Medb |
5 | Manannán (40) | Oscar (15) | Fionn | Ailill |
6 | Conall Cearnach (31) | Conchobhar (14) | Milesius | Lugh |
7 | Partholón (28) | Fergus (14) | Conchobhar | Bodb Dearg |
8 | Ailill (27) | Milesius (14) | Conaire Mór | Cormac Mac Art |
9 | Aonghus Óg(24) | Deirdre (13) | Partholón | Fionn |
10 | Dagda (23) | Diarmuid (12) | Mongán | Aonghus Óg |
11 | Bodb Dearg (21) | Oisín (12) | Ogma | Fand |
12 | Cormac Mac Art (21) | Aonghus Óg (11) | Midir | Nuada |
13 | Lugh (21) | Conaire Mór (11) | Eber | Midir |
14 | Nuada (2) | Étain (11) | Ronán | Conall Cearnach |
15 | Naoise (19) | Ethlinn (11) | Bodb Dearg | Conchobhar |
16 | Fergus (18) | Manannán(11) | Dubhthach | Balor |
17 | Conaire Mór (17) | Niall Noíghiallach (11) | Fiacha | Conaire Mór |
18 | Milesius (17) | Ogma (11) | Diarmuid | Dechtiré |
19 | Ogma (17) | Art (1) | Aedh Dubh | Mórrígán |
20 | Tuireann (17) | Balor (10) | Aonghus Óg | Bécuma |

### v2

@@ 3.1 Core network properties
| 1 | Cúchulainn (70) | Cúchulainn (30) | Cúchulainn | Manannán |
| 2 | Fionn (69) | Fionn (19) | Manannán | Cúchulainn |
| 3 | Conchobhar (58) | Lugh (17) | Dagda | Dagda |
| 4 | Medb (41) | Naoise (16) | Lugh | Medb |
| 5 | Manannán (40) | Oscar (15) | Fionn | Ailill |
| 6 | Conall Cearnach (31) | Conchobhar (14) | Milesius | Lugh |
| 7 | Partholón (28) | Fergus (14) | Conchobhar | Bodb Dearg |
| 8 | Ailill (27) | Milesius (14) | Conaire Mór | Cormac Mac Art |
| 9 | Aonghus Óg(24) | Deirdre (13) | Partholón | Fionn |
| 10 | Dagda (23) | Diarmuid (12) | Mongán | Aonghus Óg |
| 11 | Bodb Dearg (21) | Oisín (12) | Ogma | Fand |
| 12 | Cormac Mac Art (21) | Aonghus Óg (11) | Midir | Nuada |
| 13 | Lugh (21) | Conaire Mór (11) | Eber | Midir |
| 14 | Nuada (2) | Étain (11) | Ronán | Conall Cearnach |
| 15 | Naoise (19) | Ethlinn (11) | Bodb Dearg | Conchobhar |
| 16 | Fergus (18) | Manannán(11) | Dubhthach | Balor |
| 17 | Conaire Mór (17) | Niall Noíghiallach (11) | Fiacha | Conaire Mór |
| 18 | Milesius (17) | Ogma (11) | Diarmuid | Dechtiré |
| 19 | Ogma (17) | Art (1) | Aedh Dubh | Mórrígán |
| 20 | Tuireann (17) | Balor (10) | Aonghus Óg | Bécuma |